OpenStudy (anonymous):

what is the volume of a cone with height 12 diameter 22

OpenStudy (saifoo.khan):

Use http://0.tqn.com/d/math/1/5/t/L/conerr.jpg

OpenStudy (experimentx):

1/2*pi*11^2*12

OpenStudy (anonymous):

1/3*22/7*121*12

OpenStudy (anonymous):

\[V = 1/3 \times \pi r^2h\]